I need help with these problems (Please!)
1) Find the vertex and focus of the parabola whose equation is 8(y-4)=(x+1)2exponent

Answer by venugopalramana(3286)

8(y-4)=(x+1)^2
the standard form of parabola of this type is
4a(y-k)=(x-h)^2
where vertex is (h,k)and focus is (h,(k+a)).
In our example we have 4a=8 or a=2
y-4=y-k...or k=4 and x+1=x-h...or ..h=-1..so
vertex is (-1,4)
focus is (-1,4+2)...or....(-1,6)